OpenStudy (jaydelv):

If sin θ = 3 over 7 and tan θ < 0, what is the value of cos θ?

Nnesha (nnesha):

 at which quadrant sin is positive

OpenStudy (jaydelv):

First

Nnesha (nnesha):

remember that CAST rule for signs

Nnesha (nnesha):

yes first and ?

OpenStudy (jaydelv):

Oh and second

Nnesha (nnesha):

yes right sin positive in 1st and 2nd quadrant given is tan < 0 is is less than 0 so at which quadrant you should draw a right triangle ?

OpenStudy (jaydelv):

second?

Nnesha (nnesha):

yes right tan is negative in 2nd quadrant <0  draw a right triangle \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]